OpenStudy (anonymous):

if 0 degrees < x<360 degrees, solve the equation sin x= -sqrt 3/2

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

that makes no sense?

OpenStudy (anonymous):

http://i.stack.imgur.com/r8uHr.gif

OpenStudy (anonymous):

sorry made a mistake 4pi/3

OpenStudy (anonymous):

but yea use the unit circle

OpenStudy (arthur_ser):

sinx=sin(-π/3) ,(0,2π) x=2kπ - (π/3) or x=2kπ + π -(π/3) x=2kπ + (2π/3) k∈Z 0<x<2π 0<2kπ - (π/3)<2π or 0<2kπ + (2π/3)<2π solve the above inequalities for k (k is an integer)

OpenStudy (anonymous):

so the answer is 150 degrees and 300 degrees?

OpenStudy (jdoe0001):

\(\bf sin(x)= -\sqrt{\frac{3}{2}}\quad \textit{taking }sin^{-1}\textit{ to both sides}\\ \quad \\ sin^{-1}[sin(x)]= sin^{-1}\left(-\sqrt{\frac{3}{2}}\right)\implies x= sin^{-1}\left(-\sqrt{\frac{3}{2}}\right)\)

OpenStudy (jdoe0001):

\(\bf \textit{recall that }sin^{-1}[sin(\theta)]=\theta\)

OpenStudy (jdoe0001):

\(\bf sin^{-1}\left(-\sqrt{\frac{3}{2}}\right)\implies \textit{the angle(s) whose sine is }-\sqrt{\frac{3}{2}}\)

OpenStudy (jdoe0001):

so check your unit circle

OpenStudy (anonymous):

okay so then its 120 and 240? in degrees

OpenStudy (jdoe0001):

240 is in the 3rd quadrant, sine is negative there, yes 120 is in the 2nd quadrant, sine is positive there, no dice

OpenStudy (jdoe0001):

notice, is a negative \(\Large -\sqrt{\frac{3}{2}}\)

OpenStudy (anonymous):

soo then it has to be 300 then

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

thank you

OpenStudy (jdoe0001):

lemme recheck something

OpenStudy (anonymous):

ok

OpenStudy (jdoe0001):

ohh... I see.. nevermind... you meant ... to write I gather \(\bf sin(x)= -\cfrac{\sqrt{3}}{2}\quad \textit{taking }sin^{-1}\textit{ to both sides}\\ \quad \\ sin^{-1}[sin(x)]= sin^{-1}\left(-\cfrac{\sqrt{3}}{2}\right)\implies x= sin^{-1}\left(-\cfrac{\sqrt{3}}{2}\right)\\ \quad \\ x= \begin{cases} 240^o\\ \bf 300^o \end{cases}\) just making the distinction, that \(\bf -\cfrac{\sqrt{3}}{2}\ne -\sqrt{\cfrac{3}{2}}\)

OpenStudy (anonymous):

actually the signs "<" are suppose to be greater that or equal to for both so wold it still be 240 and 300?

OpenStudy (jdoe0001):

0 < x < 360 just means 0 < x <--- "x" is greater than 0 x < 360 <---- "x" is lesser than 360 well, 240 is greater than 0 and less than 360 300 is also greater than 0 and less than 360

OpenStudy (jdoe0001):

they're not BELOW 0 or equals it, they're not ABOVE 360 or equals it

OpenStudy (anonymous):

240 degrees

OpenStudy (anonymous):

ok so then it is 240 and 300

OpenStudy (jdoe0001):

yes

OpenStudy (arthur_ser):